# Stability of Non-Hyperbolic Sets

### v1

@@ 4.1 The combinatorics
Use coordinates centered at $z^{(2)}$ and let $m_{5}=(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(2)})$ be the preimage of $z^{(3)}$ under $N$ iterates, i.e. $F^{N}_{t,a}(\Delta x,\Delta y)=z^{(3)}$. Then $0\leq\Delta y\leq L\left(\lambda^{\theta}\mu\right)^{n}$. Let $v=\left(\begin{matrix}v_{1}\\ v_{2}\end{matrix}\right)$ be the tangent vector at $m_{5}$ to $W^{u}_{\text{loc}}(z^{(2)})$. We use Lemma 8 and we obtain that
$|v_{2}|=\left[2Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\right]|v_{1}|.$ (4.10)
By Lemma 1 we have that ${\partial^{2}Y}/{\partial y^{2}}$ is bounded away from zero in a neighbourhood of $c$. As a consequence, by Lemma 5 we have
$DF^{N}_{t,a}(m_{5})=\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),$
where $A$, $B$, $C$ are matrices and $D$ is a number. Observe that, from Lemma 1, there exists a uniform constant $d^{\prime}>0$ such that
$D=d^{\prime}\left(m_{5,y}-1\right)+O(m_{5,x}),$
and by Lemma 7, $m_{5,x}=O\left(\lambda^{n}\right)$. By using again the Lemma 7 and the bound on $m_{5,x}$, by taking $\epsilon_{1}$ small enough, we get that $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ where $d>0$ is a uniform constant. Because $F^{N}_{t,a}(m_{5})$ is the minimum of the curve $W^{u}_{\text{loc}}(z^{(3)})$, then
$Cv_{1}+Dv_{2}=0.$ (4.11)
Use coordinates $\left(\Delta x^{\prime},\Delta y^{\prime}\right)$ along $W^{u}_{\text{loc}}\left(z^{(1)}\right)$ centered around $z^{(1)}$ such that $F^{n}\left(\Delta x^{\prime},\Delta y^{\prime}\right)=\left(\Delta x,\Delta y\right)$. Then, from (4.7), $\Delta y^{\prime}=O\left(\lambda^{\theta n}\right)$ and by the quadratic behavior, $|\Delta x^{\prime}|=O\left(\lambda^{\frac{\theta}{2}n}\right)$. This implies that
$|\Delta x|=O\left(\lambda^{\left(1+\frac{\theta}{2}\right)n}\right).$ (4.12)
By (4.10), (4.11), Lemma 8, the fact that $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ and that $C$ is a uniform constant, we have
$2dQ_{2}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{1}{\lambda}\right)^{n}\right)=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$
Moreover by (4.12), $|\Delta x|\left(\frac{1}{\lambda}\right)^{n}=O\left(\lambda^{\frac{\theta}{2}n}\right)$ and because $\lambda^{\frac{\theta}{2}n}<1$ we obtain
$|\Delta x|=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.13)
Consider the projection of the curve $W^{u}_{\text{loc}}(z^{(2)})$ on the the $(x,y)$-plane. The curvature of this projected curve is proportional to the curvature of the curve $W^{u}_{\text{loc}}(z^{(2)})$. Because $F^{n}$ is linear, we can estimate the curvature of the projection of the curve $W^{u}_{\text{loc}}(z^{(2)})$ to the $(x,y)$-plane by calculating how the linear map, which preserves the $(x,y)$-plane, changes the curvature of $W^{u}_{\text{loc}}(z^{(1)})$ on the $(x,y)$-plane. We denote the $x$-coordinate by $x$. We calculate now the second derivatives using again Lemma 8 and the fact that the curve $W^{u}_{\text{loc}}(z^{(2)})$ is in $n$ ${\mathcal{C}}^{2}$} exponentially close to the projected curve to the $xy$-plane. We get
$\frac{d^{2}y}{dx^{2}}\geq\frac{3}{2}Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}+O\left(|\Delta x|\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\geq Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n},$ (4.14)
where we used (4.13). Observe that $W^{u}_{\text{loc}}(m_{5})$ is a graph over an $x$-direction with second derivative at $m_{5}$ given by (4.14). By (4.10) and (4.13), the slope satisfies
$\frac{dy}{dx}=O\left(\left(\frac{1}{\lambda^{\theta}\mu}\right)^{n}\right).$ (4.15)
Consider now the same curve as a graph over the $y$-axis. Then, by (4.14), (4.15) and using $d^{2}x/dy^{2}=-1/\left(dy/dx\right)^{3}d^{2}y/dx^{2}$, we have
$\frac{d^{2}x}{dy^{2}}\geq C\left(\frac{\mu^{4}}{\lambda^{2-3\theta}}\right)^{n}.$
The map $F^{N}_{t,a}$ will preserve this order of curvature. As final remark observe that, by (2.16), ${\mu^{4}}/{\lambda^{2-3\theta}}=\left({\mu}/{\lambda^{2-\theta}}\right){\mu^{3}}{\lambda^{2\theta}}>1$. ∎
@@
We are now ready to chose $n_{0}$. Let
$S_{n}=\left\{(x,y)\in[-2,2]^{m}|y\geq W_{n}(x)\right\}.$
Observe that $z\in S_{n}\setminus S_{n-1}$. Fix $t$ and let $n_{0}$ be the maximal $n\in{\mathbb{N}}$ such that
$W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))\subset S_{n-n_{0}}.$ (4.16)
Observe that such a choice of $n_{0}$ gives information on the position of $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ with respect to the pull-backs of the stable manifold, see Figure 5. A more precise estimate on the size of $n_{0}$ comes from the following lemma. Let $\alpha={\log\left(\lambda^{2\theta}\mu^{3}\right)}/{\log\mu}$. By (2.17), $\alpha\in\left(0,{1}\right)$.
Lemma 11.
The integer $n_{0}$ satisfies the following:
$n_{0}=[n\alpha]+O(1).$
@@
Fiz $t$ as in the definition of $n_{0}$. From Lemma 9, (2.8) and the fact that $z\in S_{n}\setminus S_{n-1}$, there exist uniform constants $K_{1},K_{2}>0$ such that
$z_{y}^{(3)}(t,a_{n}(t))=K_{1}\left(\frac{1}{\mu^{n}}+\left(\lambda^{\theta}\mu\right)^{2n}\right).$
Moreover the definition of $n_{0}$ implies that
@@
$\frac{1}{K}\leq\frac{1}{\mu^{n_{0}}}\left(1+\left(\lambda^{2\theta}\mu^{3}\right)^{n}\right)\leq K.$ (4.17)
The lemma follows from (2.16). ∎
Remark 10.
Observe that by combining (4.17) and (2.16), one gets
$\lambda^{2\theta}\mu^{2}<\frac{1}{\mu^{\alpha}}.$ (4.18)
This simplifies the formulas in Lemmas 7 and 9.
@@ 4.2 The secondary tangencies
In this subsection we construct secondary tangencies along the curve $a_{n}$. We start by giving the formal definition of secondary tangency.
Definition 5.
@@
We take a vector at a point $m_{1}\in W^{u}_{\text{\rm loc}}\left(c\right)$ and calculate its variation along its orbit at crucial moments, step by step, until we get a vector at the point $F^{3N+\theta n+n}(m_{1})\in W^{u}_{\text{\rm loc}}(z^{(3)})$. The steps will be defined precisely in the proof of the proposition. The reader can refer to Figure 4.
A crucial application of Proposition 2 is the speed of $W^{u}_{\text{\rm loc}}(z^{(3)})$ when varying the parameters along the curve $a_{n}$. The dominating terms in the estimates for the partial derivatives will cancel in the calculation of this speed; see (4.48). This is the reason why we cannot suppress more terms in the Taylor development. This cancellation is also crucial in the proof of Proposition 4.
Proposition 2.
For $n\geq 1$ large enough and for $(t,a)\in\mathcal{B}_{n}$,
$D(C-E)\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]\right\}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)$
@@
$\frac{DT}{\mu^{n_{0}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}/2}}\right\}+O\left(\frac{1}{\mu^{n_{0}}}\right),$
@@
$D(C-E)\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D(C-E)\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]\right\}$
$+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\leq\frac{\partial z_{y}^{(3)}}{\partial a}\leq$
$DT\mu^{n-n_{0}/2}+1+\frac{DT}{\mu^{n_{0}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}/2}}\right\}+O\left(\frac{1}{\mu^{n_{0}}}\right).$
Moreover $\lambda^{\theta}\mu^{2}>1$, CE,D,T>0C-E,D,T>0 and they converge to a non zero limit as $n\to\infty$.
In particular, when $a=a_{n}$,
$\frac{\partial z_{y}^{(3)}}{\partial t}=DC\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right\}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right),$
@@
$\frac{\partial z_{y}^{(3)}}{\partial a}$ $=$ $DC\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+DC\left(\lambda^{\theta}\mu\right)^{n}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]\right\}$
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$
@@
Choose $t_{1}\in[-t_{0},t_{0}]$. Observe that, by Lemma 5, the definition of unfolding, by using coordinates $(x,y,t,a)$ centered around $(0,1,t_{1},0)$ in the domain and coordinates centered around $F_{t_{1},0}^{N}(0,1)$ we get the following expression for $F_{t,a}^{N}(x,y)$,
$F_{t,a}^{N}(x,y)$ $=$ $\left(\begin{matrix}Ax+By+tE+aF\\ Cx+Qy^{2}+a\left[1+x\varphi_{x}+y^{2}\varphi_{y}\right]\end{matrix}\right),$ (4.19)
where the uniformly bounded coefficients satisfy
    $A(x,y,t,a)$ is a $(n-1)\times(n-1)$ matrix valued ${\mathcal{C}}^{3}$} function,
    $B(x,y,t,a),E(x,y,t,a),F(x,y,t,a)$ are n1n-1 vector valued ${\mathcal{C}}^{3}$} functions,
    $C(x,y,t,a)$ is a $1\times(n-1)$ matrix valued ${\mathcal{C}}^{3}$} function,
    $Q(x,y,t,a),\varphi_{x}(x,y,t,a),\varphi_{y}(x,y,t,a)$ are ${\mathcal{C}}^{3}$} functions.
Moreover $B,Q\neq 0$. The $x$ component of $F_{t,a}^{N}(x,y)$ is a general expression and the fact that $B\neq 0$ follows from $(f5)$. For the $y$ component of $F_{t,a}^{N}(x,y)$ observe that the linear terms in $t$ and $y$ are absent because in $a=0$, the point $F^{N}_{t,0}(0,1)$ is a non-degenerate homoclinic tangency, see Remark 6. The $a$-dependence of the second component of $F^{N}_{t,a}$ follows from Remark 6.

### v3

@@ 4.1 The combinatorics
In order to be able to use Lemma 2.2.12, use coordinates $(\Delta x_{c},\Delta y_{c})\in W^{u}_{\text{loc}}(z^{(2)})$ centered at $c$ and coordinates $(\Delta x_{z},\Delta y_{z})\in W^{u}_{\text{loc}}(z^{(3)})$ centered at $z$. By Lemma 2.2.12, we have
$\Delta x_{z}$ $=$ $F^{N}\left(\Delta x_{c}\right)=A\Delta x_{c}+B\Delta y_{c}+O\left(\Delta x_{c}^{2}+\Delta x_{c}\Delta y_{c}+\Delta y_{c}^{2}\right)$ (4.1.35)
$\Delta y_{z}$ $=$ $F^{N}\left(\Delta y_{c}\right)=C\Delta x_{c}+Q\Delta y_{c}^{2}+O\left(\Delta x_{c}^{2}+\Delta x_{c}\Delta y_{c}+\Delta y_{c}^{3}\right).$ (4.1.36)
For the aim of proving the lemma we are interested in coordinates $(\Delta x_{z^{(2)}},\Delta y_{z^{(2)}})\in W^{u}_{\text{loc}}(z^{(2)})$ centered at $z^{(2)}$ and coordinates $(\Delta x_{z^{(3)}},\Delta y_{z^{(3)}})\in W^{u}_{\text{loc}}(z^{(3)})$ centered at $z^{(3)}$. By construction, by Lemma 4.1.21, Lemma 4.1.23 and Lemma 4.1.22 we have,
$\Delta x_{z}$ $=$ $\Delta x_{z^{(3)}}+1-{z_{x}^{(3)}}=\Delta x_{z^{(3)}}+O\left(H\right),$
$\Delta y_{z}$ $=$ $\Delta y_{z^{(3)}}+{z_{y}^{(3)}}-z_{y}=\Delta x_{z^{(3)}}+O\left(H^{2}\right),$
$\Delta x_{c}$ $=$ $\Delta x_{z^{(2)}}+O\left(\lambda^{n}\right),$
$\Delta y_{c}$ $=$ $\Delta y_{z^{(2)}}+H=\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}+H+O\left(\left(\lambda^{\frac{3}{2}\theta}\mu\right)^{n}\right).$
Using the equalities above in (4.1.35), (4.1.36) and after several order cancellations, we get
$\Delta x_{z^{(3)}}$ $=$ $H\left[B+O(1)\right]+A\Delta x_{z^{(2)}}\left[1+O(H)\right]+\frac{Q}{B}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}\left[1+O(H)\right]$ (4.1.37)
$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{3}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}\Delta x_{z^{(2)}}^{4}\right)=f\left(\Delta x_{z^{(2)}}\right),$ (4.1.38)
$\Delta y_{z^{(3)}}$ $=$ $H^{2}\left[Q+O(1)\right]+C\Delta x_{z^{(2)}}\left[1+O(H)\right]+2H\frac{Q^{2}}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}\left[1+O(H)\right]$ (4.1.39)
$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{3}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}H\Delta x_{z^{(2)}}^{4}+\left(\frac{\mu^{3}}{\lambda^{6}}\right)^{n}\Delta x_{z^{(2)}}^{6}\right)=g\left(\Delta x_{z^{(2)}}\right).$ (4.1.40)
By Lemma 4.1.22,
$\Delta y_{z^{(3)}}=\frac{C_{2}}{2}\Delta x_{z^{(3)}}^{2}+O\left(\Delta x_{z^{(3)}}^{3}\right)$ (4.1.41)
with $C_{2}\neq 0$. We need then to calculate the coefficient $C_{2}$. Observe that
$C_{2}=\frac{d^{2}\Delta y_{z^{(3)}}}{d\Delta x_{z^{(3)}}^{2}}$ $=$ $g^{\prime\prime}\left(\Delta x_{z^{(2)}}\right)\left(\frac{d\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}}\right)^{2}+g^{\prime}\left(\Delta x_{z^{(2)}}\right)\frac{d^{2}\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}^{2}}$ (4.1.42)
$=$ $g^{\prime\prime}\left(\Delta x_{z^{(2)}}\right)\left(\frac{d\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}}\right)^{2},$
where we used that $z^{(3)}$ is the lowest point of the $W^{u}_{\text{\rm loc}}(z^{(3)})$ and hence $g^{\prime}\left(\Delta x_{z^{(2)}}\right)=0$. By (4.1.39),
$g^{\prime}=\frac{d\Delta y_{z^{(3)}}}{d\Delta x_{z^{(2)}}}$ $=$ $C\left[1+O(H)\right]+4H\frac{Q^{2}}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O(H)\right]$
$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}H\Delta x_{z^{(2)}}^{3}+\left(\frac{\mu^{3}}{\lambda^{6}}\right)^{n}\Delta x_{z^{(2)}}^{5}\right),$
and
$g^{\prime\prime}=\frac{d^{2}\Delta y_{z^{(3)}}}{d\Delta x_{z^{(2)}}^{2}}$ $=$ $4H\frac{Q^{2}}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\left[1+O(H)\right]$
$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}H\Delta x_{z^{(2)}}^{2}+\left(\frac{\mu^{3}}{\lambda^{6}}\right)^{n}\Delta x_{z^{(2)}}^{4}\right).$
By the previous inequality and by (4.1.31) after cleaning the order terms we get
$g^{\prime\prime}\left(\Delta x_{z^{(2)}}\right)=4H\frac{Q^{2}}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\left[1+O(H)\right].$ (4.1.43)
Observe that
$\frac{d\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}}=\frac{1}{f^{\prime}\left(\Delta x_{z^{(2)}}\right)}$ (4.1.44)
and by (4.1.37),
$f^{\prime}=\frac{d\Delta x_{z^{(3)}}}{d\Delta x_{z^{(2)}}}$ $=$ $A\left[1+O(H)\right]+2\frac{Q}{B}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O(H)\right]$
$+$ $O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}+\left(\frac{\mu^{2}}{\lambda^{4}}\right)^{n}\Delta x_{z^{(2)}}^{3}\right).$
In particular, by (4.1.33)
$f^{\prime}\left(\Delta x_{z^{(2)}}\right)=-\frac{CB}{2Q}\frac{1}{H}\left[1+O(H)\right],$
and by (4.1.44)
$\frac{d\Delta x_{z^{(2)}}}{d\Delta x_{z^{(3)}}}=-\frac{2Q}{CB}{H}\left[1+O(H)\right].$ (4.1.45)
Use (4.1.43) and (4.1.44) in (4.1.42) to get
$C_{2}$ $=$ $\frac{16Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\left[1+O(H)\right],$
and by (4.1.41),
$\Delta y_{z^{(3)}}=\frac{8Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(3)}}^{2}+O\left(\Delta x_{z^{(3)}}^{3}\right).$
@@
For all $(t,a)\in\mathcal{B}_{n}$, let $n_{0}$ be the maximal $n\in{\mathbb{N}}$ such that
$z^{(3)}\subset\left\{(x,y)\in[-2,2]^{2}|y\geq W_{n-n_{0}}(x)\right\}.$ (4.1.46)
Observe that the funtion $(t,a)\mapsto n_{0}$ gives information on the position of $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ with respect to the pull-backs of the stable manifold, see Figure 5. A more precise estimate on the size of $n_{0}$ comes from the following lemma. Let
$\alpha=\frac{\log\left(\lambda^{2\theta}\mu^{3}\right)}{\log\mu}$ (4.1.47)
and recall the definition of $n_{0}^{*}$ in (4.0.1). By (2.3.2), $\alpha\in\left(0,{1}\right)$.
Lemma 4.1.48.
For every $(t,a_{n}(t))\in\mathcal{B}_{n}$, the integer $n_{0}$ satisfies the following:
$n_{0}=[n\alpha]+O(1)>n_{0}^{*}.$
In particular there exists a uniform positive constant $K$ such tha
$\frac{1}{K}\left(\lambda^{2\theta}\mu^{3}\right)^{n}\leq\mu^{n_{0}}\leq K\left(\lambda^{2\theta}\mu^{3}\right)^{n}.$ (4.1.49)
@@
Let $(t,a_{n}(t))\in\mathcal{B}_{n}$. From Lemma 4.1.23, (2.2.2) and the fact that $W_{n+1}(1)<z_{y}<W_{n}(1)$, there exists a uniform constant $K_{1}>0$ such that
$\frac{1}{K_{1}}\left(\frac{1}{\mu^{n}}+\left(\lambda^{\theta}\mu\right)^{2n}\right)\leq z_{y}^{(3)}\leq K_{1}\left(\frac{1}{\mu^{n}}+\left(\lambda^{\theta}\mu\right)^{2n}\right).$
Moreover the definition of $n_{0}$ implies the existence of a uniform constant $K_{2}>0$ such that
@@
$\frac{1}{K}\leq\frac{1}{\mu^{n_{0}}}\left(1+\left(\lambda^{2\theta}\mu^{3}\right)^{n}\right)\leq K.$
The lemma follows from (2.3.1). ∎
Remark 4.1.50.
Observe that by combining (4.1.47) and (2.3.1), one gets
$\lambda^{2\theta}\mu^{2}<\frac{1}{\mu^{\alpha}}.$ (4.1.51)
@@ 4.2 The secondary tangencies
In this subsection we locate secondary tangencies along the curve $a_{n}$. We start by giving the formal definition of secondary tangency.
Definition 4.2.1.
@@
We take a point $m_{1}\in W^{u}_{\text{\rm loc}}\left(c\right)$ and consider its orbit at crucial moments, step by step, until we get to the point $F^{3N+\theta n+n}(m_{1})\in W^{u}_{\text{\rm loc}}(z^{(3)})$. In particular we are interested in the derivative of these points with respect to $m_{1},t$ and $a$. The crucial moments are the ones defined by the combinatorics as presented in the previous section. The reader can refer to Figure 4.
The reader might be surprised by the smaller order terms we need to keep track of in Proposition 4.2.2. Unfortunately the leading terms will cancel and the speed of $W^{u}_{\text{\rm loc}}(z^{(3)})$ when varying the parameters along the curve $a_{n}$ will be determined by smaller order terms, see (4.2.33). Smaller order terms also turned out to be crucial in the proof of Proposition 5.3.1.
Proposition 4.2.2.
For $n\geq 1$ large enough and $(t,a)\in\mathcal{B}_{n}$,
$DS_{C-E}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\}$
@@
$\frac{DT}{\mu^{n_{0}^{*}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}^{*}/2}}+O\left(\frac{1}{\mu^{n_{0}^{*}/2}}\right)\right\},$
@@
$DS_{C-E}\left\{\mu^{n}+\frac{1}{DS_{C_{E}}}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\}$
$\leq\frac{\partial z_{y}^{(3)}}{\partial a}\leq$
$DT\mu^{n-n_{0}^{*}/2}+1+\frac{DT}{\mu^{n_{0}^{*}/2}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}^{*}/2}}+O\left(\frac{1}{\mu^{n_{0}^{*}/2}}\right)\right\},$
where $S_{C-E}=(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$. Moreover C,CE,DC,C-E,D are positive smooth functions in $t,a$ and depending also on $n$. These functions converge to a non zero function in $t$ as $n\to\infty$. The constant $T$ is the same as in the definition of $\mathcal{B}_{n}$. In particular, when $a=a_{n}(t)$,
$\frac{\partial z_{y}^{(3)}}{\partial t}=DS_{C}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\},$
@@
$\frac{\partial z_{y}^{(3)}}{\partial a}$ $=$ $DS_{C}\left\{\mu^{n}+\frac{1}{DS_{C}}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\},$
where $S_{C}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$.
@@
Choose $t_{1}\in(-t_{0},t_{0})$. By Lemma 2.2.12 and by Remark 2.1.15, we get the following expression for $F_{t,a}^{N}(x,y)$, in coordinates $(x,y,t,a)$ centered around $(0,1,t_{1},0)=(c,t_{1},0)$ and coordinates centered around $z=F_{t_{1},0}^{N}(0,1)$,
$F_{t,a}^{N}(x,y)$ $=$ $\left(\begin{matrix}Ax+By+tE+aF\\ Cx+Qy^{2}+a\left[1+x\varphi_{x}+y^{2}\varphi_{y}\right]\end{matrix}\right),$ (4.2.3)
where $A(x,y,t,a),B(x,y,t,a),C(x,y,t,a),E(x,y,t,a),F(x,y,t,a),\varphi_{x}(x,y,t,a)$ are ${\mathcal{C}}^{3}$} functions and $Q(x,y,t,a),\varphi_{y}(x,y,t,a)$ are ${\mathcal{C}}^{2}$} function over $B,Q\neq 0$. The fact that $B\neq 0$ follows from $(f4)$. For the $y$ component of $F_{t,a}^{N}(x,y)$ observe that the linear terms in $t$ and $y$ are absent because in $a=0$, the point $F^{N}_{t,0}(0,1)$ is a non-degenerate homoclinic tangency, see Remark 2.1.15. The $a$-dependence of the second component of $F^{N}_{t,a}$ follows from Remark 2.1.15.